OpenStudy (anonymous):

Solve for x. 8 over x equals x over 11. 2 radical 22, 22 radical 2, 44 or 88?

OpenStudy (anonymous):

8 over x equals x over 11. 8/x = x/11 8(11)=x^2 x=sqrt(88) x=2sqrt22 ans or 2 radical22

OpenStudy (anonymous):

Thank you.